General Informations

Status-Key

Status/green.gif

Resolved

Status/orange.gif

Still working but with some errors

Status/red.gif

Pending

Maintenance itet-stor service

STATUS: Status/green.gif

The server behind the itet-stor service will be updated from Debian Wheezy to Debian Jessie in the morning of 2015-10-29. The service will be temporarily unavailable.


Maintenance work on D-ITET's central IT infrastructure

STATUS: Status/green.gif

2015-09-12: 17:30
All services are back online
2015-09-12: 10:00-18:00
To keep our systems up to date with the newest software and security releases, we need to update our servers on a regular basis. For this reason we are going to install latest Oracle patches on our main servers. The servers will be rebooted during this maintenance

To prevent data corruption/loss please do the following:

More details:


TerminalServer vega7 is running instable

STATUS: Status/green.gif

2015-07-22: 12:05
  • The MS TerminalServer vega7 disconnects users unexpectedly. Users claim to get a temporary profile. Please do not connect to vega7 until those issues are fixed.

2015-07-22: 18:30
  • Configuration error found. Repair is in progress.


No Access to Inboxes

STATUS: Status/green.gif

2015-07-22: 20:43
  • Emails sent to users create bounces because the mailserver can not access the users inbox.
2015-07-22: 22:13
  • A permission problem was the reason for this problem. We are still investigating what caused this permission problem. The Mailserver should work now as intended.


Major Problems in network connectivity to D-ITET Servers

STATUS: Status/green.gif

2015-05-14: 08:00
  • We are experiencing major networking problems to and from D-ITET servers causing disturbance to logins, webservice and mail. We have informed ID networking team about the issues.
2015-05-14: 17:20
  • Network is stabilizing again and services should soon be back to normal.
2015-05-14: 18:50
  • Network issues started again.
2015-05-14: 21:20
  • ID networking team is working in the issue.
2015-05-14: 21:45
  • ID networking team has isolated a suspicious possibly hacked server. We have shut down the system and will investigate the issue. Network is going back to normal.
2015-05-15: 13:55
  • We have analyzed the affected system and found the break in which was used to mount a denial of service. Firewalls at the D-ITET network boundaries detected the anomal traffic and started to drop SYN packages. During this time there was only a ~0.5% success probability to create a new TCP connection from D-ITET network to outside, causing the disturbance to all services.


Overall System Downtime Due To Maintenance Work

STATUS: Status/green.gif

2015-05-09: 11:55
  • All services are back online
2015-05-09: 9:00 - 12:00
  • Due to some major modifications in the user configuration causing a system wide impact we are going to shutdown ALL services.

    To prevent data corruption or loss please do the following:

  • save and close all open files
  • close all running applications
  • logout from all ITET systems
  • do not establish any connection from outside

    NOT logged off users will be removed without warning! This includes the corresponding processes as well!


VEGA7 Terminal Server with heavy load

STATUS: Status/green.gif

2015-03-24: 14:30
  • VEGA7 was in an unstable state. Only an enforced reboot was able to recover all crashed services.
2015-03-24: 14:30
  • VEGA7 is on heavy load, login is not possible anymore. We are investigating the problem.


Problems on core server drwho

STATUS: Status/green.gif

2015-02-27: 21:30
  • Systems recovering and going back to normal state.
2015-02-27: 17:00 - 22:00
  • Planned maintenance downtime:

    ITET's core server DRWHO experiences major problems with its disk controller and its associated disks. Affected hardware needs to be replaced by Oracle support. Due to this maintenance work we are going to shutdown DRWHO during this time slot.

    Primarily LINUX desktops and compute servers are concerned, i.e all Linux desktops and compute servers are down as well.
    To prevent data loss please do the following:

    • save all open files
    • close all running applications
    • logout from all ITET LINUX systems
    • shutdown your personal ISG.EE managed LINUX Desktop
    Managed WINDOWS computers and email services are not affected by this downtime.
2015-02-13: 09:15
All services should now be available again. The maintenance downtime was much longer than planned. Please apologize for any inconveniences that might have caused.
2015-02-13: 08:30
The problems seem not to be resolved. The server is online but some services are not running yet.
2015-02-13: 08:00
The firmware update took longer than planned. The server should be back online in the next minutes.
2015-02-12: 11:00
The firmware needs to be updated. This will be done on Friday, 13th of February, between 06:30 and 07:30 hours.
2015-02-09: 14:30
The server is back online, we are still investigating the issue.
2015-02-09: 13:15
Affected are all diskless clients and services which depend on drwho (e.g. the email.ee.ethz.ch website).
2015-02-09: 12:55
The D-ITET server drwho.ee.ethz.ch is currently not reachable.


www.ee.ethz.ch and www.werkstatt.ee.ethz.ch - Websites down

STATUS: Status/green.gif

2015-01-26: 18:53

Informatikdienste resolved the problems and http://www.ee.ethz.ch and http://www.werkstatt.ee.ethz.ch are now again online.

2015-01-26: 14:12
ID annouces, that the downtime could last until 18:30.
2015-01-26: 09:30
Informed ID Servicedesk about the issue.


lists.ee.ethz.ch - Composing has been temporarily disabled

STATUS: Status/green.gif

2015-01-20: 22:05
We applied the security updates and released again the composing area.
2015-01-20: 11:15
Composing of messages has been temporarily disabled due to a security flaw. People using the composing tool are able to view and send any file on the filesystem. Composing will be re-enabled after the package maintainers have updated sympa.


SMTP Email Sending: Sophos Problems

STATUS: Status/green.gif

2015-01-15: 17:45
After the last virus-signature update of the Sophos antivirus software, we had a virus data checksum error. This was the reason for the delays. After fixing this issue, everything went to normal state.
2015-01-15: 16:00
The sending and receiving of emails takes too much time and the mailserver is on high load. We are working on a solution.


Server oenone NFS problems

STATUS: Status/green.gif

2015-01-12: 21:37

We see NFS related problems on server oenone again. Affected were homes of all users from AUT, BIWI, IKT and VAW. This issue also affects related webpages and services.

2015-01-13: 07:20
The system is again recovered and back online.


CategoryEDUC

Status/Archive/2015 (last edited 2016-01-06 13:10:19 by alders)